ADNAN MUHAMMED

Grade 12,

Q.Maximum number of electrons are present in.option are A.2.24 lit.of SO2 at NTP. B.0.2 molesof NH3 . C.1.5 gm atoms of oxygen. D.2 gram atoms of sulphur

Q.Maximum number of electrons are present in.option are A.2.24 lit.of SO2 at NTP.  B.0.2 molesof NH3 . C.1.5 gm atoms of oxygen. D.2 gram atoms of sulphur

Grade:11

2 Answers

Khimraj
5 years ago
electrons  = mole*NA*electron in one molecule
  1. (2.24/22.4)*NA*32
  2. 0.2*NA*10
  3. (1.5/16)*NA*16
  4. (2/32)*NA*32
So correct answer is (a)
Khimraj
5 years ago
 
electrons  = mole*NA*electron in one molecule
2.24 lit.of SO2 at NTP = (2.24/22.4)*NA*32
  1. 0.2 molesof NH3 = 0.2*NA*10
  2. 1.5 gm atoms of oxygen= (1.5/16)*NA*16
  3. 2 gram atoms of sulphur = (2/32)*NA*32
So correct answer is (a)